The LSU ATIC crew (Gould, Granger, Guzik, Isbert, Smith, Stewart)
arrived on-site at the Willy Field LDB facility on Wednesday October 24.
We were able to unpack both the instrument
and support equipment C-containers and, within a few hours, had a
minimum ATIC system powered up and collecting muon calibration data.
Work toward resolving some
payload building issues on Wednesday was cut short when we needed to
evacuate the site due to worsening weather.
We were able to return to the payload building on Thursday (October 25)
morning. The overnight muon data was downloaded and analyzed.
Results from this analysis showed that all hodoscope and BGO calorimeter
detector channels were fully functional and no major repairs /
adjustments are required. The remainder of the day was spent keeping
out of the way of CSBF and RPSC personnel as they brought Payload
building 1 up to full operational status. I need to very strongly
commend both Dave Sullivan and Phil Austin for their heroic and
effective effort in rallying the McMurdo resources required to solve
that building issues.
On Friday, October 26, we will able to use the entire payload building
and complete the initial ATIC setup. The ATIC control and monitoring
systems were established in their permanent location for the season on
the mezzanine. In addition, the ATIC instrument was moved to the
pressure vessel ring during the morning and to the lower part of the
external structure by 3 pm in the afternoon. This is, in fact, a new
record for ATIC assembly. The experiment is currently powered up and
will be collecting muon calibration data throughout the night.
Following analysis of this data, we expect that ATIC will be ready for
installation of the SIP tomorrow.
Tonight is the annual McMurdo Halloween party.
The Flight accommodations on a C-17
in the External
The ATIC 4experiment in
the lower 1/2 of the external frame.
Track the ATIC experiment